OpenStudy (anonymous):

Is this right? Identify the correct slope and y-intercept of the equation 2x + 3y = 6 slope = -2/3; y-intercept = (0, 2)

OpenStudy (anonymous):

i think so 2x+3y=y 3y=6 y=6 ,(0,6) 2x+3y=6 3y=-2x+6 y=-2/3x+2 slope =-2/3

OpenStudy (anonymous):

yea i did the same thing thanks